OpenStudy (jacky90):

The time it takes to cover the distance between two cities by car varies inversely with the speed of the car. The trip takes 7 hours for a car moving at 30 mph How long does the trip take for a car moving at 42 mph?

OpenStudy (shadowlegendx):

So instead of varying directly like y = kx, it will vary inversely like y = k/x

OpenStudy (shadowlegendx):

Danno do your magic

OpenStudy (518nad):

30 miles per hour, so in 7 hours thats 30*7 = 210 miles travelled in total 42 miles per hour, to travel 210 miles, you see how many 42 miles are in 210, each of them takes one hour 210/42=#hours it takes to travel moving at 42 mph

OpenStudy (518nad):

210/42 = 5, 5 hours

OpenStudy (jacky90):

Thank you so much! @518nad